# Smurfs: Leveraging Multiple Proficiency Agents with Context-Efficiency for Tool Planning

### v3

@@ Abstract
The emergence of large language models (LLMs) has opened up unprecedented possibilities for automating complex tasks that are often comparable to human performance. Despite their capabilities, LLMs still encounter difficulties in completing tasks that require high levels of accuracy and complexity due to their inherent limitations in handling multifaceted problems single-handedly. This paper introduces ‘Smurfs’, a cutting-edge multi-agent framework designed to revolutionize the application of LLMs. By seamlessly transforming a conventional LLM into a synergistic multi-agent ensemble, Smurfs can enhance the model’s ability to solve complex tasks at no additional cost. This is achieved through innovative prompting strategies that allocate distinct roles within the model, thereby facilitating collaboration among specialized agents and forming an intelligent multi-agent system. Our empirical investigation on both open-ended task of StableToolBench and closed-ended task on HotpotQA showcases Smurfs’ superior capability in intricate tool utilization scenarios. Notably, Smurfs outmatches all the baseline methods in both experiments, setting new state-of-the-art performance. Furthermore, through comprehensive ablation studies, we dissect the contribution of the core components of the multi-agent framework to its overall efficacy. This not only verifies the effectiveness of the framework, but also sets a route for future exploration of multi-agent LLM systems. Our code is available at https://github.com/FreedomIntelligence/Smurfs/.
Smurfs: Leveraging Multiple Proficiency Agents with Context-Efficiency for Tool Planning
Junzhi Chen [^1] These authors contributed equally to this work., Juhao Liang [^1] These authors contributed equally to this work., Benyou Wang🖂 Shenzhen Research Institute of Big Data The Chinese University of Hong Kong, Shenzhen wangbenyou@cuhk.edu.cn
@@ 1 Introduction
Tool manipulation has traditionally been seen as a distinctive human characteristic, dating back approximately 2.5 million years Oakley and Museum (1972); Ambrose (2001). For large language models (LLMs), access to external tools can equip them with broader capabilities beyond their fixed language modeling knowledge. For example, the search engine API empowers ChatGPT to access real-time information Zhao et al. (2023). However, LLMs still encounter several challenges when using multiple tools to solve tasks. These challenges include effective solution planning and adaptability to new tools. Hao et al. (2024); Guu et al. (2020); Qin et al. (2024).
This paper addresses the critical research problem of enhancing the problem-solving capabilities of LLMs through the adoption of a plug-and-play multi-agent system (MAS) framework Dorri et al. (2018); Van der Hoek and Wooldridge (2008). We posit that a MAS approach can significantly augment the efficacy of LLMs in handling tasks that require a high degree of precision, adaptability, and comprehensive knowledge integration.
@@
To this end, we introduce ‘Smurfs’ an innovative MAS framework inspired by the collaborative and versatile nature of its namesake cartoon characters. The proposed framework is based on the principle: synergistic collaboration among specialized agents can overcome the limitations faced by individual LLMs. Each agent within the Smurfs framework is designed to perform specific sub-tasks, facilitating a more nuanced and effective approach to complex problem-solving. Our research delves into the architectural design, coordination mechanisms, and the operational dynamics of integrating specialized agents into a cohesive system. The effectiveness of Smurfs is validated through both open-ended and closed-ended tool planning benchmark experiments Guo et al. (2024); Yang et al. (2018), where the proposed MAS system consistently outperform baseline methods on both benchmarks. An ablation study followed by a case study further investigates the underlying reasons for this effectiveness. These results not only establish a new state-of-the-art in the field but also offer concrete evidence of the multi-agent approach’s efficacy in enhancing LLM capabilities.
@@
    We introduce a novel plug-and-play MAS framework to enhance the tool planning capabilities of LLMs. Experiments demonstrate the effectiveness of this approach, which is also more cost-efficient compared to existing tool planning methods.
@@
    Ablation studies further reveal the underlying reasons for the effectiveness of the MAS framework, providing valuable insights for future research.
2 Motivation
2.1 Multi-Tool Planning
To augment LLMs to do multi-tool planning for solving complex problems, previous work has seen numerous attempts. Chain-of-Thought  Wei et al. (2023) was the first to propose the method of thought and answer chain reasoning. ReACT Yao et al. (2022) further introduced the thought-action-observation format for tool chain reasoning, leading to the development of various multi-tool planning methods Chen et al. (2023a); Xu et al. (2023); Shinn et al. (2023). The latest work, DFSDT Qin et al. (2024), was proposed to address the inherent limitations of CoT and ReACT: error propagation and limited exploration. Deep First Search Decision Tree, denoted as DFSDT, is powerful in addressing multi-tool planning problems. Its core concept involves employing a depth-first search (DFS) approach for multi-tool planning (for more details, see Appendix A). When a tool fails or is deemed inadequate for solving the current problem, DFSDT backtracks to the previous solution state and attempts to resolve the issue using a different tool. However, several limitations were identified with the mechanism of DFSDT: (1) instability of the rollback mechanism, (2) redundant context, and (3) premature termination. The following sections will introduce these limitations in detail.
2.1.1 Instability of the Rollback Mechanism
@@ 2.1.3 Premature Termination
2.2 Multi Agent System
@@
Table 2: Comparison of related works.
To address the limitations inherent in DFSDT and to further enhance LLM’s multi-tool planning capabilities, multi-agent system (MAS) has emerged as a natural solution. Inspired by human social division of labor and cooperation, MAS aim to enable AI agents to accomplish more complex tasks more effectively and efficiently through the division of labor and collaboration. Previous works Song et al. (2023); Liu et al. (2023); Chen et al. (2023b); Yin et al. (2024); Qiao et al. (2024) has leveraged MAS to achieve this goal. Table 2 shows the difference between them. Based on those works, we further design the MAS named Smurfs to address issues with DFSDT. By dividing tasks among different agents, each agent can focus on a specific part of the DFSDT task, accessing only the necessary history as context during task execution, which effectively addresses the issue of redundant context. The redesign of the rollback mechanism to incorporate memory and tool list rollback mechanisms addresses the instability of the rollback mechanism. Drawing on the concept of least-to-most prompting Zhou et al. (2023), the original problem is first decomposed into sub-problems for macro-level planning. Subsequently, DFSDT is used to solve each sub-problem at the micro-level, with macro-level planning guiding the micro-level planning, thereby resolving the issue of premature termination.
@@
3 Smurfs: A framework with multiple agents
@@ 3.1 Framework Overview
Figure 1 illustrates the entire workflow for the Smurfs framework. Initially, the Planning Agent identifies the user’s complex request and breaks it down into manageable sub-tasks. Executor Agents are then tasked with collecting task specific information, utilizing access to external tools. Answer Agent compiles the findings into a cohesive response, which is subsequently verified by the Verifier Agent to ensure accuracy and relevance. Each agent focus on its own task and only use the relevant part of the conversation history to reduce the Redundant Context. This process exemplifies the framework’s capability to efficiently handle complex queries by leveraging the specialized roles of multiple agents, thereby ensuring both the precision of task execution and the quality of the output. In the following sections, the system mechanism and functions of each agent will be detailed. More details of memory system can be seen at Appendix B.
@@ 3.4 Subtask Solving Process
Where q is sub-problem from the Planning Agent, r is response from the Executor Agent, M is the local memory (or global memory if max retry reaches). As the ‘lost-in-the-middle’ theory described in section 2.1, retaining all information may not always be beneficial, particularly in cases where the solution path is challenging to discern. Therefore, the primary role of the Answer Agent is to succinctly summarize the generated answers and tool responses to maintain the memory efficiency.
@@ Results
We conducted a detailed analysis of the token costs associated with each tool planning method for the tasks, a critical evaluation aspect for multihop reasoning tasks. As shown in Table 1 (detailed in Appendix E), the average token costs per question and API request are evaluated for ReACT, DFSDT, and Smurfs on StableToolBench. The analysis reveals that DFSDT generally requires about 20,000 tokens per question, encompassing both prompt and completion tokens. This is nearly three times the token cost compared to ReACT and twice as much as Smurfs. Despite this higher token cost, DFSDT does not demonstrate commensurate effectiveness improvements over other methods. These findings underscore the cost-efficiency of the proposed MAS framework, Smurfs, which not only reduces token expenditure in solving multihop planning tasks but also delivers outstanding performance in evaluations.
@@ 4.2 Closed-ended Task: HotpotQA
Following the settings in  Qiao et al. (2024), we use open-source Llama-2 models Touvron et al. (2023) and Mistral-7B Jiang et al. (2023) as the backbones of each agent to evaluate the performance of Smurfs. The evaluation metrics is $\texttt{reward}\in[0,1]$, defined as the F1 score grading between the prediction and ground-truth answer. For more details about the experiment, see Appendix C.
@@ Results
Observations indicate that the performance of LLama-2-13b-chat on smurfs-related tasks is suboptimal, likely due to its limited capabilities in tool arguments generation. Specifically, the primary issue identified is that, when the Executor agent successfully selects relevant tool, it tends to produce hallucination arguments that can’t be used by the tools. This indicates that LLama-2-13b-chat may need further training for usage of tools. The experimental results may substantiate this viewpoint, demonstrating that the untrained methods of llama-2-13b-chat generally exhibit significantly lower accuracy compared to the trained methods. Nevertheless, Smurfs achieves the second highest accuracy among the untrained methods, only slightly behind reflexion, which still attests to Smurfs’ capability.
@@ 5.1 Importance of each component in MAS
We performed an ablation study to investigate the impact of each agent in our framework. We removed each agent individually, except for the indispensable Executor Agent, and compared the results to the complete framework. Table 5 shows that the Planning Agent is the most crucial component, followed by the Verifier Agent, with the Answer Agent being the least important.
(1) Verifier Agent Removal: Without verification, the framework uses a general depth-first search, leading to increased computational demand and more tool invocations.
(2) Answer Agent Removal: Removing this agent means the Executor Agent’s answers won’t be summarized, risking the ’lost-in-the-middle’ problem due to lengthy tool responses. As shown in the results, a more intelligent model, GPT-4 Turbo, can mitigate the negative impact of the Answer Agent’s removal. We believe this is because the more powerful model can leverage more information effectively.
(3) Planning Agent Removal: Removing the Planning Agent affects the global path-searching strategy. Models with Smurfs may show reduced performance without preliminary planning, as seen in current frameworks like ReACT and DFSDT. The results demonstrate that the impact of removing the Planning Agent is significant, as it directly influences the multihop reasoning ability of the MAS.
@@ 6 Conclusion
In this study, we present a novel MAS framework, ‘Smurfs’, tailored to enhance the planning and reasoning capabilities of LLMs in handling complex tasks that involve lengthy contexts and tools. We conduct experiments on the multi-step tool usage benchmark, StableToolBench and HotpotQA, and the results demonstrate the overall effectiveness and efficiency of the Smurfs framework compared to baseline methods.
In conclusion, this research contributes to the expanding field of study focused on enhancing LLM capabilities, particularly for multi-step tool usage tasks. It emphasizes the importance of task decomposition, preliminary planning, and efficient verification for improving task execution performance. For future work, we believe incorporating more dedicated and specific roles within the system may further enhance effectiveness and efficiency, based on the ‘Smurfs principle’: synergistic collaboration among specialized agents can overcome the limitations faced by individual LLMs.
7 Limitations
@@
Due to computational constraints, our experiments did not include larger and more diverse types of LLMs.
@@
Appendix A Details of DFSDT
See Figure 5.
Figure 5: A comparison of DFSDT and conventional CoT or ReACT during model reasoning (left) Qin et al. (2024). We show part of the solution path annotation process using GPT-4 (right).
Appendix B Details of the Smurfs
See Figure 6 for executor working process and Figure 4 for memory and tool library of Smurfs.

### v4

# Smurfs: Multi-Agent System using Context-Efficient DFSDT for Tool Planning

@@ Abstract
Teaching large language models (LLMs) to use tools for solving complex problems can grant them human-like reasoning abilities. ReAct and its variants are popular frameworks for tool use in both single-agent and multi-agent systems. To address issues like error propagation and limited exploration in ReAct, the Deep First Search Decision Tree (DFSDT) was proposed, but it faces challenges such as rollback instability, redundant context, and premature termination in single-agent settings. We introduce "Smurfs," a novel multi-agent system (MAS) that enhances DFSDT with a modular, context-efficient, and training-free design. Smurfs surpasses baseline methods in both the open-ended StableToolBench and the closed-ended HotpotQA tasks, reducing token usage by 60.9% compared to DFSDT and enabling Mistral-7b to perform on par with GPT-4-DFSDT. Extensive ablation studies confirm the effectiveness of Smurfs’ core components, offering valuable insights for the construction and interpretation of MAS, and paving the way for future exploration. We release the code at https://github.com/FreedomIntelligence/Smurfs.
Smurfs: Multi-Agent System using Context-Efficient DFSDT for Tool Planning
Junzhi Chen [^1] These authors contributed equally to this work., Juhao Liang [^1] These authors contributed equally to this work., Benyou Wang🖂 The Chinese University of Hong Kong, Shenzhen Shenzhen Research Institute of Big Data wangbenyou@cuhk.edu.cn
@@ 1 Introduction
The ability to manipulate tools for complex tasks has long been considered a distinctive characteristic of humans Oakley and Museum (1972); Ambrose (2001). Can we extend this ability to today’s large language models (LLMs), enabling them to utilize multiple tools to perform complex tasks beyond their inherent capabilities? If LLMs can use external tools to access knowledge or execute tasks beyond their fixed language modeling capabilities, we can shift the focus of LLM training towards enhancing their reasoning and tool-use skills. This shift would allow tools to supplement what LLMs should know or execute, thereby improving the parameter efficiency of the LLMs.
@@
In previous multi-agent systems for tool planning, methods like Chain-of-Thought (CoT) Wei et al. (2023), ReAct Yao et al. (2022), and the more advanced DFSDT Qin et al. (2024) have been proposed to enhance LLMs’ ability to handle complex multi-step tasks. However, these approaches face notable limitations. ReAct has trouble in handling error propagation and limited exploration. DFSDT incorporates a rollback mechanism and depth-first search approach to address limitations of ReAct, but it suffers from instability when the base model struggles with long-context reasoning. It also introduces inefficiencies due to redundant context handling and risks premature termination when solving multi-step problems. These challenges highlight the need for further innovation to better manage context and reasoning complexity in multi-tool planning systems.
In this paper, we introduce ‘Smurfs’, an innovative multi-agent system (MAS) framework inspired by the collaborative and versatile nature of its namesake cartoon characters. The proposed framework leverages enhanced DFSDT to perform complex tool planning tasks without the need for additional training. The effectiveness of Smurfs is demonstrated through both open-ended and closed-ended tool planning benchmark experiments Guo et al. (2024); Yang et al. (2018), where it consistently outperforms baseline methods. An ablation study, followed by a case study, further investigates the reasons behind this effectiveness. These results establish a new state-of-the-art in the field and provide concrete evidence of the advantages of a multi-agent approach in enhancing LLM capabilities.
@@
    We introduce a highly modular, context-efficient, and training-free MAS framework that utilizes an enhanced DFSDT to improve the tool planning capabilities of LLMs. Experiments demonstrate the effectiveness of this approach, which also proves to be more cost-efficient compared to existing methods.
@@
    Through ablation studies, we uncover the underlying factors contributing to the effectiveness of the MAS framework, offering valuable insights for future research.
2 Related Work
To augment LLMs to do multi-tool planning for solving complex problems, previous work has seen numerous attempts. Chain-of-Thought  Wei et al. (2023) is the first to propose the method of thought and answer chain reasoning. ReAct Yao et al. (2022) further introduce the thought-action-observation format for tool chain reasoning, leading to the development of various multi-tool planning methods Chen et al. (2023a); Xu et al. (2023); Shinn et al. (2023). The latest work, Deep First Search Decision Tree (DFSDT) Qin et al. (2024), is proposed to address the inherent limitations of CoT and ReACT: (1) Error Propagation: Error occurs at early stage of planning will result in wrong answer in the end, but it can only be identified until reaching the end of the planning chain (2) Limited Exploration: Single solution chain can’t explore the planning space completely.
DFSDT is powerful in addressing multi-tool planning problems. Its core concept involves employing a depth-first search (DFS) approach for multi-tool planning and backtracks whenever an LLM think the solving process has entered a wrong state (for more details, see Appendix B). When a tool fails or is deemed inadequate for solving the current problem, DFSDT backtracks to the previous solution state and attempts to resolve the issue using another solution plan. However, three limitations are identified with the mechanism of DFSDT:
Limit I: Instability of the Rollback Mechanism
@@ 2 Related Work
In conclusion, DFSDT relies highly on the base model’s reasoning ability, especially long context reasoning ability to make roll back decision, termination decision and tool choice decision at the same time, which is a very difficult task even with the most powerful LLM like GPT4 Yuan et al. (2024).
Multi-Agent System
To address the limitations inherent in DFSDT and to further enhance LLM’s multi-tool planning capabilities, MAS has emerged as a natural solution. Inspired by human social division of labor and cooperation, MAS aim to enable AI agents to accomplish more complex tasks through the division of labor and collaboration. By decomposing the task of DFSDT to multiple agents and giving them only the information they need, we can enable LLMs to use DFSDT more effectively and more efficiently.
@@
Table 2: Comparison of tool use systems.
Previous work on multi-agent system mainly focus on coding and society simulation area Hong et al. (2023); Qian et al. (2024); Park et al. (2023); Li et al. (2023); Wu et al. (2023). For tool-use scenario, most multi-agent systems still use the ReAct style reasoning Qiao et al. (2024); Chen et al. (2023b); Yuan et al. (2024); Liu et al. (2023); Song et al. (2023); Yin et al. (2024); Xu et al. (2023), only using multi-agent discussion and revision to increase reasoning quality, which still inherits the limitation of ReAct. Therefore, this paper aims to construct a novel MAS framework to address the aforementioned limitations. Table 2 shows the difference between different tool-use systems. More detailed discussion can be seen in Appendix A.
Token Compression
Token compression refers to compressing tokens fed into LLMs while preserving inference performance, thus reducing computational overhead and mitigating the constraints imposed by long context limits. Previous works Mu et al. (2024); Fu et al. (2024) have explored token compression techniques, with a focus on compressing token embeddings. In contrast, Smurfs achieves context compression by filtering the input context for each tool planning process. Table 6 shows the detailed difference between different compression techniques.
In the future, these compression methods can be used together in multi-tool planning scenarios to achieve more efficient token compression. Each agent’s system prompt could be compressed using gist tokens Mu et al. (2024), tool descriptions could be managed with CAMPHOR’s compression approach Fu et al. (2024), and the context input by each agent could be compressed using Smurfs.
@@
3 Smurfs: MAS with Context Efficient DFSDT
@@ 3.1 Framework Overview
Figure 1 illustrates the entire workflow for the Smurfs framework. Initially, the Planning Agent identifies the user’s complex query and breaks it down into manageable sub-tasks. Executor Agents are then tasked with collecting task specific information, utilizing access to external tools. Answer Agent compiles the findings into a cohesive response, which is subsequently verified by the Verifier Agent to ensure accuracy and relevance.
By dividing tasks among different agents, each agent can focus on a specific part of the task, accessing only the necessary history as context during task execution, which effectively addresses the issue of redundant context. The redesign of the rollback mechanism to incorporate memory and tool list rollback mechanisms addresses the instability of the rollback mechanism. Drawing on the concept of least-to-most prompting Zhou et al. (2023), the original problem is first decomposed into sub-problems for macro-level planning. Subsequently, Smurfs is used to solve each sub-problem at the micro-level, with macro-level planning guiding the micro-level planning, thereby resolving the issue of premature termination.
In the rest of this section, the mechanism of the system and the functions of each agent will be detailed. More details of memory system can be seen at Appendix C.
@@ 3.2 Agent Components
In this section, we introduce the two core components of the Smurfs system:
@@ 3.4 Subtask Solving Process
Where q is sub-problem from the Planning Agent, r is response from the Executor Agent, M is the local memory (or global memory if max retry reaches). As the ‘lost-in-the-middle’ theory described in section 1, retaining all information may not always be beneficial, particularly in cases where the solution path is challenging to discern. Therefore, the primary role of the Answer Agent is to succinctly summarize the generated answers and tool responses to maintain the memory efficiency.
@@ Results
We conducted a detailed analysis of the token costs associated with each tool planning method for the tasks, a critical evaluation aspect for multihop reasoning tasks. As shown in Table 1 (detailed in Appendix G), the average token costs per question and API request are evaluated for ReACT, DFSDT, and Smurfs on StableToolBench. The analysis reveals that DFSDT generally requires about 20,000 tokens per question, encompassing both prompt and completion tokens. This is nearly three times the token cost compared to ReACT and twice as much as Smurfs. Despite this higher token cost, DFSDT does not demonstrate commensurate effectiveness improvements over other methods. These findings underscore the cost-efficiency of the proposed MAS framework, Smurfs, which not only reduces token expenditure in solving multihop planning tasks but also delivers outstanding performance in evaluations.
@@ 4.2 Closed-ended Task: HotpotQA
Following the settings in  Qiao et al. (2024), we use open-source Llama-2 models Touvron et al. (2023) and Mistral-7B Jiang et al. (2023) as the backbones of each agent to evaluate the performance of Smurfs. The evaluation metrics is $\texttt{reward}\in[0,1]$, defined as the F1 score grading between the prediction and ground-truth answer. For more details about the experiment, see Appendix D.
@@ Results
Observations indicate that the performance of LLama-2-13b-chat on smurfs-related tasks is suboptimal, likely due to its limited capabilities in tool arguments generation. Specifically, the primary issue identified is that, when the Executor agent successfully selects relevant tool, it tends to produce hallucination arguments that can’t be used by the tools. This indicates that LLama-2-13b-chat may need further training for usage of tools. The experiment results may substantiate this viewpoint, demonstrating that the untrained methods of llama-2-13b-chat generally exhibit significantly lower accuracy compared to the trained methods. To provide further insight into the model’s performance in the tool planning process, we manually categorized the types of errors made by Smurfs on the hotpotQA hard dataset in appendix E.
@@ 5.1 Importance of each component in MAS
We performed an ablation study to investigate the impact of each agent in our framework. We removed each agent individually, except for the indispensable Executor Agent, and compared the results to the complete framework.
Settings
(1) Verifier Agent Removal: Without Verifier Agent, the framework uses a generatl DFSDT in each subtask, i.e. including a finish tool into the tool library and deciding whether to stop at tool choice step. (2) Answer Agent Removal: Without Answer Agent, the framework uses full tool response instead of the summary of tool response in its memory. (3) Planning Agent Removal: Without Planning Agent, the framework uses Verifier Agent to decide whether the task is completed. (4) Planning & Verifier Agent Removal: Without Planning and Verifier Agent, the only difference between Smurfs and DFSDT will be the Answer Agent, i.e. including a finish tool in the tool library, deciding whether to stop at tool choice step and using summary of tool response in its memory.
Results
Table 5 shows the experiment result, highlighting several key insights regarding the impact of different components in the multi-agent system.
First, the removal of any component generally reduces the win rate, highlighting each component’s significance. Second, performance degradation trends remain consistent: Eliminating the Answer Agent has minimal impact, whereas removing the Planning Agent causes the greatest decline. However, this does not imply the relative importance of these agents but rather suggests that the Verifier-Planning Agent workflow is more robust than the Verifier-Answer Agent workflow.
Third, the performance impact varies by model capability. For GPT-3.5, the removal of the Answer Agent reduces both the pass and the win rates, while for GPT-4, it preserves the win rate and improves the pass rate, probably due to the superior long-text reasoning ability of GPT-4, which processes additional tool response more effectively. Removal of the Planning Agent significantly reduces the pass rate of GPT-3.5 but has a minimal effect on GPT-4, suggesting that the GPT-4 Verifier Agent is more robust. These findings may indicate that more powerful models can compensate for agent removal, sometimes even benefiting from additional context, while weaker models benefit more from complex workflow and compressed context.
We also noticed an interesting phenomenon: when only the Planning Agent is removed, the system experiences a major decline in performance, while removing the Verifier Agent in addition to the Planning Agent improves performance again. This may indicate that the Verifier Agent’s ability to pause tasks or provide next-step guidance is suboptimal and requires further training.
In conclusion, our findings offer some insight into the relationship between model capabilities and multi-agent system performance. The optimal number of agents and workflow may vary depending on the capacity of the model. We propose the hypothesis that weaker models benefit more from complex multi-agent systems and context segmentation, while stronger models perform better with comprehensive context and simpler agent systems. To validate this hypothesis, more extensive ablation studies are needed under a wider range of models and constraints to explore the influence of different context structures and reasoning workflows. We leave this investigation for future work.
@@ 6 Conclusion
In this paper, we propose Smurfs, an innovative MAS framework designed to enhance the tool-planning capabilities of LLMs without requiring additional training. Through extensive experiments on both open-ended and closed-ended tool planning benchmarks, Smurfs demonstrate its effectiveness by consistently outperforming baseline methods. Ablation study further provides deeper insights into the impact of each agent in our framework. Based on the ablation study, we propose a hypothesis that can be further tested through more comprehensive experiments in future research: weaker models benefit more from complex multi-agent systems and context segmentation, while stronger models perform better with comprehensive context and simpler agent systems. The findings not only advance the state-of-the-art in multi-tool planning systems but also highlight the potential of modular, training-free frameworks for LLMs in various practical applications.
Looking forward, future research could focus on exploring Smurfs’ use in new domains, such as facilitating the synthesis of high-quality multi-tool planning data and enhancing the base model’s reasoning and tool-use abilities, further advancing the field of adaptive AI systems.
Limitations
@@
Due to computational constraints, our experiments did not include larger and more diverse types of LLMs. We believe this would not affect the main observations of this paper.
@@
Acknowledgements
This work was supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (grant number ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023.
Appendix A Comparison between Smurfs and existing MAS
In conclusion, Smurfs stands out compared to existing MAS in followig terms:
Flexibility
Smurfs utilize enhanced version of DFSDT, which intergrades global and iterative planning, while existing MAS mainly use ReAct and iterative planning. This make Smurfs more flexible in the planning process.
Adaptability
Smurfs realize superior performance on multiple different benchmarks, proving the adaptability of Smurfs. Existing MAS like RestGPT, Lumos are tailored for specific downstream task and need additional training to be used in other scenario.
Learning Efficiency
Smurfs do not need training, thus have the highest learning efficiency among existing MAS. It only need query and tool documentation from the user.
To further illustrate these points, we conducted a detailed comparison between Smurfs and two well-known Multi-Agent Systems, highlighting their differences and the adjustments required when learning out-of-box tasks.
CAMEL
CAMEL Li et al. (2023) is a communicative agent framework. It uses role-play technique and inception-prompting technique to achieve autonomous cooperation between agents. CAMEL does not natively support tool use settings. CAMEL is considered to perform poorly when generalizing to new tasks (e.g., on the MATH dataset Wu et al. (2023)).
Autogen
Autogen Wu et al. (2023) is not designed as an agent framework for any specific task scenario. Instead, it provides a multi-agent conversation framework that allows users to customize agent characteristics and complete tasks through discussions among different agents. Autogen addresses various problems not by employing a uniform workflow, but by allowing users to design customized agents and workflows flexibly based on their tasks.
Smurfs
Smurfs is designed as a unified workflow specifically for complex multi-tool planning scenarios. For different tasks, Smurfs only needs to adjust the few-shot examples in the agent prompts and provide documentation for the tools applicable to the task, allowing Smurfs to generalize to other task scenarios. Smurfs was initially designed for the Stable-toolbench, which itself encompasses various types of tasks and has access to over 16,000 plugins. HotpotQA was subsequently introduced to evaluate the performance of Smurfs on closed-end tasks. When migrating from Stabletoolbench to HotpotQA, only the few-shot examples in the planning agent prompts were modified, along with the provision of plugin documentation for HotpotQA. The rest of the system continued using the same unified framework as in Stabletoolbench. Experiment results demonstrate that our untrained out-of-box unified framework achieves and even surpasses performance of agent systems specifically trained on HotpotQA, such as Autoact and Fireact, showcasing Smurfs’ flexibility in generalization.
Compression Method Applicable Scenario Compressed Object Implementation
Gist Tokens Mu et al. (2024) General scenarios Frequently used system prompts Training LLM to compress system prompts, reducing token usage.
CAMPHOR Fu et al. (2024) Tool use scenarios Tool descriptions Adopts a similar approach to gist tokens, compressing each tool description into a single token, thus reducing token cost.
Smurfs Complex multi-tool planning scenarios Input context in tool planning Operates on multi-tool planning workflows, compressing the context needed for each tool planning process.
Table 6: Comparison of Token Compression Methods
Appendix B Details of DFSDT
DFSDT Qin et al. (2024) gives control to the model to stop and rollback the solution trajectory by using Finish tool, thus addressing limitations of ReACT. Finish tool has two parameters give answer: model thinks the task is finished and decide to give answer and give up and restart: model thinks current trajectory can’t lead to correct answer and decide to rollback.
Appendix C Details of the Smurfs
Executor Agent Details
As illustrated in Figure 5, given a sub-problem p, Executor Agent first thinks about what to do this time, generates thought $\gamma$ according to p, local memory M, hint h from the Verifier Agent and tool list at the current step $\tau$. Then it will choose action $\alpha$ to complete the sub-problem using p, $\gamma$ and $\tau$. After that, parameters of $\alpha$ are generated using p, local memory M and tool description of the action $D[\alpha]$. Tool is then invoked to complete the task.
Memory Details
As illustrated in Figure 4, there are four kinds of memory in Smurfs. Local memory stores thought-answer pairs of the current solution trajectory, while global memory stores all history solution trajectory (including those that is backtracked). Tool list only stores available tools’ name and its usage description, while tool doc stores all detailed information about the tools including parameters details. Through using different kinds of memory under different circumstances, Smurfs can use DFSDT in a context efficient way.
Restart Mechanism
Every time Smurfs generate an intermediate output, a format checker is used to check whether the output is of the expected format. If not, Smurfs will retry the same step until reach retry limit or generate correct format output. This mechanism is used in addition to the rollback mechanism to handle the situation where the system can generate correct content but fail to follow the output format.